xolile has more moneythan yvonne. if xolile gave yvonne R200, they would have the same amount of money. if yvonne gave R220, xolile would have twice as much as yvonne. let x be the amount of money xolile has and y be the amount of money yvonne has. what is the suitable equations for this??
Answer by macston(5194)

x-200=y+200
x=y+400
STOP HERE: The equations above are sufficient to solve the problem. Solution continues below.
x+220=2(y-220) Substitute for x.
y+400+220=2y-440
0=y-1060
1060=y ANSWER 1: Yvonne has R1060
x=y+400=1460 ANSWER 2: Xolile has R1460